Find (a) the reduced form. (b) the numerical value of π‘Œπ‘’ , and (c) the effect on the multiplier if a proportional income tax (t) is incorporated into the model. π‘Œ = 𝐢 + 𝐼, 𝐢 = 𝐢𝑂 + π‘π‘Œ 𝑑 , 𝑇 = 𝑇𝑂 + π‘‘π‘Œ, π‘Œ 𝑑 = π‘Œ βˆ’ 𝑇 where 𝐼 = 𝐼𝑂 = 30, 𝐢𝑂 = 85, 𝑏 = 0.75, 𝑑 = 0.2, and 𝑇𝑂 = 20


1
Expert's answer
2021-03-11T18:24:50-0500

Solution:

a.). The reduced form:

Y = AD

First derive structural equation:

Ye = C + I

Ye = CO + bYd + IO

Ye = CO + b (Ye – (TO + tYe)) + IO

Formulate reduced form:

Ye = CO + b (Ye – (TO + t)) Ye + IO

Ye - b (1 – (TO + t)) Ye = CO + IO

[1 – b (1 – (TO + t)] Ye = CO + IO

Solve for Ye:


Ye ="\frac{C_{O} + I_{O}}{1 – b (1 – (T_{O} + t)) }"

Β 

b.). The numerical value of Ye:

Ye = "\frac{C_{O} + I_{O}}{1 – b (1 – (T_{O} + t)) }"


Ye = "\frac{85 + 30}{1 – 0.75 (1 – (20 + 0.2)) }"


Ye = "\frac{115}{15.40}" = 7.47


Ye = 7.47

Β 

d.). If a proportional income tax (t) is incorporated into the model, the multiplier effect will be low, that is the multiplier will decrease.